# An AI-powered Bayesian generative modeling approach for causal inference in observational studies

### v2

@@ 3.5 Binary Treatment Experiments
Furthermore, we evaluate whether the CausalBGM framework can learn a more effective low-dimensional representation compared to CausalEGM and sufficient dimension reduction (SDR). It is important to note that all SDR-based methods for causal inference rely on linear SDR, which is inherently restrictive and may fail to capture nonlinear relationships in complex datasets. To assess this, we conducted a comprehensive comparison of CausalBGM with SDRcausal under experimental settings that either satisfied or violated the SDR assumption. SDRcausal implements several variants proposed in the original study (Ghosh et al., 2021), and for fairness, we always report the best-performing result. CausalBGM demonstrated significant improvements over SDRcausal in both experimental settings, with substantial advantages in nonlinear datasets where the linear SDR approach was unable to model the underlying complexity effectively (see Appendix E). These results highlight the capability of CausalBGM to overcome the limitations of linear assumptions and better capture intricate relationships in high-dimensional data.
@@ 3.7 Effect of Initialization
The EGM initialization strategy plays an important role in ensuring the superior performance of CausalBGM. To evaluate the contribution of the EGM initialization strategy, we conducted a series of experiments comparing it with the traditional Xavier uniform initializer (Glorot and Bengio, 2010) across three simulation datasets and one semi-synthetic dataset under the continuous treatment setting. The results, summarized in Table 3, demonstrate that EGM initialization significantly enhances the performance of CausalBGM in terms of both RMSE and MAPE. Quantitatively, EGM initialization consistently reduced RMSE across all datasets and improved MAPE in three out of four datasets. For instance, in the Lee et al. dataset, EGM initialization achieved remarkable reductions in RMSE and MAPE by 93.4% and 80.1%, respectively. Similarly, in the Imbens et al. and Sun et al. datasets, EGM initialization substantially improved performance, with RMSE reductions of 70.5% and 78.4%, respectively. Even in the Twins dataset, where the impact on MAPE was marginal, EGM initialization still achieved a noticeable RMSE improvement of 50.0%.
@@ 3.8 Scalability
Scalability has become a critical requirement in causal inference, particularly for modern applications involving increasingly large and complex datasets. To evaluate the scalability of CausalBGM, we conducted comprehensive experiments examining its ability to handle datasets with a high number of covariates and large sample sizes. Our results demonstrate that CausalBGM is capable of processing datasets with over 50,000 covariates and more than 1 million samples with a reasonable computational resource, achieving reliable and consistent performance (See Appendix F). In contrast, many competing methods struggled or failed to handle datasets of this magnitude, highlighting the superior scalability of CausalBGM. We also showed the running time of CausalBGM under different sample sizes (See Appendix G). These findings underscore the practicality of CausalBGM in addressing the computational demands of large-scale causal inference problems in real-world applications.
@@ 4 Discussion
In this article, we introduced CausalBGM, a powerful and scalable Bayesian generative modeling framework for causal inference, particularly excelling in observational studies with high-dimensional covariates and large-scale datasets. By combining the principles of three domains: AI, Bayesian inference, and causal inference, CausalBGM provides a flexible and robust approach to analyze the complex causal relationships among variables while ensuring statistical rigor.
@@
One of the most significant contributions of CausalBGM is its ability to estimate posterior intervals for individual treatment effects (ITEs), an area that has been largely overlooked by existing causal inference methods. CausalBGM adopts a Bayesian framework and uses an iterative algorithm to infer individual-level posterior distributions of latent features. This innovation allows for the construction of well-calibrated posterior intervals at the individual level, offering a new perspective on understanding causal effects that is critical for applications requiring personalized decision-making. Additionally, the scalability of CausalBGM lies in the design of the iterative updating algorithm, which only requires a mini-batch of samples or a single sample for each step. Such scalability, combined with its robust statistical foundations, makes CausalBGM a practical and powerful tool for addressing the demands of modern applications in genomics, healthcare, and social sciences.
@@
Despite these strengths, certain limitations provide opportunities for future improvement. First, while CausalBGM demonstrates strong empirical performance with its Bayesian foundation, further theoretical work is needed to rigorously characterize the convergence properties of CausalBGM under the proposed iterative algorithm. Second, the sensitivity of CausalBGM to parameter initialization remains poorly understood, which could limit its adaptability in scenarios where the EGM framework is less effective. Investigating the underlying causes of this sensitivity and exploring alternative initialization strategies or adaptive learning mechanisms could further enhance the robustness and versatility of the framework. Third, identifiability is a critical issue in latent variable modeling, particularly for ensuring valid causal inferences. We discuss how to address this issue in Appendix H by leveraging nonlinear Independent Component Analysis (ICA) theory. Fourth, using variational inference (VI) may lead to an underestimation of the posterior uncertainty (Murphy, 2012), we investigate the effect of VI for our model in Appendix I.
@@
In conclusion, CausalBGM provides a new perspective on developing Bayesian causal inference methods by harnessing the power of AI. Its flexibility, scalability, and strong empirical performance make it a valuable tool for a wide range of applications. By addressing both theoretical and practical challenges, future iterations of CausalBGM have the potential to further advance causal inference methodologies and broaden their impact in modern data-driven applications.

### v3

@@ 3.5 Binary Treatment Experiments
Furthermore, we evaluate whether the CausalBGM framework can learn a more effective low-dimensional representation compared to CausalEGM and sufficient dimension reduction (SDR). It is important to note that all SDR-based methods for causal inference rely on linear SDR, which is inherently restrictive and may fail to capture nonlinear relationships in complex datasets. To assess this, we conducted a comprehensive comparison of CausalBGM with SDRcausal under experimental settings that either satisfied or violated the SDR assumption. SDRcausal implements several variants proposed in the original study (sdrCausal2021), and for fairness, we always report the best-performing result. CausalBGM demonstrated significant improvements over SDRcausal in both experimental settings, with substantial advantages in nonlinear datasets where the linear SDR approach was unable to model the underlying complexity effectively (see Appendix E). These results highlight the capability of CausalBGM to overcome the limitations of linear assumptions and better capture intricate relationships in high-dimensional data.
@@ 3.7 Effect of Initialization
The EGM initialization strategy plays an important role in ensuring the superior performance of CausalBGM. To evaluate the contribution of the EGM initialization strategy, we conducted a series of experiments comparing it with the traditional Xavier uniform initializer (pmlr-v9-glorot10a) across three simulation datasets and one semi-synthetic dataset under the continuous treatment setting. The results, summarized in Table 3, demonstrate that EGM initialization significantly enhances the performance of CausalBGM in terms of both RMSE and MAPE. Quantitatively, EGM initialization consistently reduced RMSE across all datasets and improved MAPE in three out of four datasets. For instance, in the Lee et al. dataset, EGM initialization achieved remarkable reductions in RMSE and MAPE by 93.4% and 80.1%, respectively. Similarly, in the Imbens et al. and Sun et al. datasets, EGM initialization substantially improved performance, with RMSE reductions of 70.5% and 78.4%, respectively. Even in the Twins dataset, where the impact on MAPE was marginal, EGM initialization still achieved a noticeable RMSE improvement of 50.0%.
@@ 3.8 Scalability
Scalability has become a critical requirement in causal inference, particularly for modern applications involving increasingly large and complex datasets. To evaluate the scalability of CausalBGM, we conducted comprehensive experiments examining its ability to handle datasets with (i) increasing the number of covariates and (ii) increasing the sample size (see Appendix F).
We first fixed number of covariates and increased the sample size using Imbens et al dataset. Only OLS, CausalEGM and CausalBGM can scale to million-sample dataset (e.g., $N=10^{6}$) while other approaches failed due to memory or time limitations. We then fixed the sample size and increased the number of covariantes from $p=50$ to p=10,000p=10{,}000. In this setting, several competing methods (e.g., DML) failed to run as $p$ increased while CausalBGM demonstrated consistently strong performance across the entire range of covariate dimensions. We further showed the computational time of CausalBGM by varying sample size over orders of magnitude from N=1, [^0] N=1{,}000 to 1,000,0001{,}000{,}000 (See Appendix G).
@@ 4 Discussion
In this article, we introduced CausalBGM, an AI-powered Bayesian generative modeling framework for causal inference, particularly excelling in observational studies with high-dimensional covariates and large-scale datasets. By combining deep generative modeling with Bayesian principles and causal reasoning, CausalBGM provides a flexible and robust approach for flexible covariate adjustment, individualized effect estimation, and principled uncertainty quantification.
@@
One notable contribution of CausalBGM is the ability to estimate individual-specific posterior distributions of latent features (interpretable as latent confounders), enabling posterior intervals for individual treatment effects (ITEs) in addition to point estimates. This individualized posterior representation offers a practically useful and statistically grounded way to quantify uncertainty at the individual level, which remains underdeveloped in existing causal inference methods. Additionally, CausalBGM is designed to be scalable since the iterative updating algorithm only requires a mini-batch of samples at each step. Such scalability, combined with its robust statistical foundations, makes CausalBGM a practical and powerful tool for addressing the demands of modern applications in genomics, healthcare, and social sciences.
Despite these strengths, several limitations and open questions remain, and we explicitly address key methodological claims to improve transparency, reproducibility, and interpretability. First, The latent variables are typically unidentifiable for latent variable modeling (LVM) since different latent representations can induce the same marginal distribution over the observed variables. We discuss this issue by leveraging recent advances in nonlinear independent component analysis (ICA) theory (see Appendix H). Specifically, identifiability can be achieved under mild conditions by introducing an auxiliary variable and conditioning the latent prior on it. Building on this idea, we developed an identifiable variant, iCausalBGM and evaluated the empirical performance compared to CausalBGM.
Second, using variational inference (VI) may lead to an underestimation of the posterior uncertainty (murphy2012machine), we implement a “full MCMC” version of CausalBGM where Metropolis–Hastings is used for sampling individual latent variables and Hamiltonian Monte Carlo is used for sampling model parameters. The empirical performance is benchmarked against the original CausalBGM (see Appendix I).
Third, CausalBGM relies on a latent-variable version of unconfoundedness, which requires relevant confounding information can be captured by observed covariates. When important covariates are missing, the impact depends on their causal role. We evaluate the performance impact on CausalBGM under different “missing-covariate” scenarios to demonstrate the robustness of our method (see Appendix J).
@@
Looking forward, several directions could further strengthen the framework. On the theory side, while CausalBGM demonstrates strong empirical performance with its Bayesian foundation, further theoretical work is needed to rigorously characterize the convergence properties of CausalBGM under the proposed iterative algorithm. On the methodological side, the sensitivity of CausalBGM to parameter initialization remains poorly understood, which could limit its adaptability in scenarios where the EGM framework is less effective. Investigating the underlying causes of this sensitivity and exploring alternative initialization strategies or adaptive learning mechanisms could further enhance the robustness and versatility of the framework.
@@
In conclusion, CausalBGM provides a new perspective on developing Bayesian causal inference methods by harnessing the power of AI. Its flexibility, scalability, and strong empirical performance make it a valuable tool for a wide range of applications in modern data science.
5 Funding
Q.L. was partially supported by NIH grant R00 HG013661. W.H.W. was partially supported by NSF-DMS 2310788.
6 Disclosure Statement
The authors report there are no competing interests to declare.